# Gyro-based Neural Single Image Deblurring

### v2

@@ 3.3 Training Strategy
Despite the gyro refinement block and gyro deblurring block, it is still challenging to train a deblurring network with erroneous gyro data. Naïvely training the network with erroneous gyro data can lead the network to ignore the gyro data and simply rely only on the blur information in the blurred image. Thus, to train our network to fully utilize input gyro data, we introduce a training strategy based on curriculum learning [2], which has shown superior performance in various learning-based approaches that deal with erroneous data [37, 39, 42].
Our training strategy initially trains the network with error-free gyro data, and gradually increases error in the training data. To this end, we first assume that for each blurred image $B$, its corresponding error-free gyro data $G_{GT}$ is given as well as its ground-truth sharp image $S_{GT}$. From $G_{GT}$, we compute a noise-free camera motion field $\mathcal{V}_{clean}$ and a noisy camera motion field $\mathcal{V}_{noisy}$. We generate $\mathcal{V}_{noisy}$ by randomly perturbing the rotational center and adding noise to the gyro data $G_{GT}$. We then compute a blended camera motion field $\mathcal{V}_{\alpha}=(1-\alpha)\cdot\mathcal{V}_{clean}+\alpha\cdot\mathcal{V}_{noisy}$ where $\alpha$ is a blending parameter. Finally, we train our network with a PSNR loss [4, 3], which is defined as:
$\mathcal{L}_{PSNR}=10\cdot\log_{10}(\|f(B,\mathcal{V}_{\alpha})-S_{GT}\|^{2}+\epsilon)$ (1)
where $f(B,\mathcal{V}_{\alpha})$ represents a deblurred output of GyroDeblurNet and $\epsilon$ is a small constant to prevent the negative infinity, which is set to $81e-$ 8 in our experiments. We initially set $\alpha$ to 0, and gradually increase it to $1$ during training. For more details including the noisy camera motion field generation and the scheduling of $\alpha$, refer to the supplementary material.
@@ 4 GyroBlur Dataset
Figure 5: Examples of the GyroBlur-Synth and GyroBlur-Real datasets. In (a), the top and bottom rows show ground-truth images and the corresponding blurry images, respectively.
@@ 4.1 GyroBlur-Synth
GyroBlur-Synth provides 14,600 and 640 synthetically blurred images of size $720\times 1280$ for the training and test sets. The dataset also provides ground-truth sharp images, ground-truth clean gyro data. To faithfully mimic real-world blurred images, the dataset covers various sources of image degradation, such as camera shakes, moving objects, noise, and saturation. The dataset also includes gyro error data to produce realistic erroneous gyro data from the clean gyro data. Fig. 5 (a) shows examples of GyroBlur-Synth.
@@
To create synthetic blurred images with moving objects that have different blur trajectories, we used a simple method. We select a single instance from the COCO dataset [27] and assume it moves at a constant speed in a blurry image. We randomly choose its position and the direction and speed of motion. We then add the object to the warped sharp images before blending them during the blur synthesis process described earlier. This method may seem simple, but our experiment shows that it helps train our model to better handle spatially-variant motion blur from moving objects.
To reflect real-world noise and saturated pixels, we also adopt the blur synthesis pipeline of RSBlur [34]. Specifically, we synthesize blur in the linear sRGB space following the aforementioned process. Then, we synthesize saturated pixels, convert the image into the camera RAW space, synthesize noise, and convert the noisy RAW image to the linear sRGB image. We refer the readers to [34] for the detailed blur synthesis process. The blur synthesis pipeline of RSBlur requires camera parameters such as a color correction matrix and a noise distribution. To this end, we use the camera parameters of a Samsung Galaxy S22 ultra-wide camera, and the noise distribution estimated from it.
The gyro error data includes gyro noise and randomly perturbed rotation centers for generating noisy camera motion fields. To obtain realistic gyro noise, we measured the gyro noise distribution from a stationary smartphone (Samsung Galaxy S22) placed on a table. Gyro noise is modeled as a Gaussian distribution for each rotation axis independently. For more details on the dataset, we refer the readers to the supplementary material.
Our synthetic dataset generation approach requires minimal device-specific information, including a small amount of gyro data sequences and parameters. This approach minimizes the need for laborious manual data collection. Specifically, we can collect a sufficient amount of gyro data by holding a smartphone in hand only for a couple of minutes, and by placing a smartphone on a table only for a few seconds. The camera parameters required for the RSBlur pipeline can also be easily obtained by a simple calibration process with a couple of images captured by a target camera. Note also that our synthetic dataset approach process does not explicitly model the relative position of the gyro sensor with respect to the camera module, which is information that is hard to access. Instead, our approach implicitly models it as part of rotational center error. Despite this, our experiments show that our approach can effectively remove blur.
@@ 4.2 GyroBlur-Real
GyroBlur-Real provides 100 real-world blurry images and their corresponding gyro data for qualitative evaluation. The dataset was collected using a Samsung Galaxy S22. The blurry images are of size $4080\times 3060$, and provided in both raw DNG and JPEG formats. The dataset also provides metadata such as timestamps of the images and gyro data, camera intrinsic parameters, exposure times and ISO parameters. Fig. 5 (b) shows examples of GyroBlur-Real.
@@ 5 Experiments
For evaluation, we implemented our network using PyTorch. We trained our network on randomly cropped $256\times 256$ image patches from the GyroBlur-Synth dataset for 300 epochs with a batch size of 16. We used the Adam optimizer [20] with $\beta_{1}=0.9$ and $\beta_{2}=0.999$. The learning rate was initially set to 0.0001 and reduced to 1e-7 using the cosine annealing scheduler [28]. We measured the computation times of all the models on a GeForce RTX 3090 GPU.
One well-known issue in using gyro data is that the timestamps of the gyro sensor and the camera are not synchronized. Fortunately, several approaches that can successfully synchronize the gyro sensor and camera have been proposed [9]. Thus, in our work, we assume that the input gyro data are already synchronized for the exposure time of the input image. GyroBlur-Real dataset was collected using an open-source Android application [1] that supports synchronization between the gyro and camera based on gyro-based clock synchronization [9].
In the remainder of this section, we present the performance evaluation of GyroDeblurNet. We evaluate the performance of our method using the test set of GyroBlur-Synth and GyroBlur-Real. Additional experiments and examples including visualization of the effect of gyro refinement and an analysis on the robustness to gyro errors are included in the supplementary material.
Table 1: Quantitative comparison on the test of GyroBlur-Synth. The inference times of the models that use gyro data include both gyro data embedding (e.g., camera motion field construction) and deblurring processes.
  Gyro Models PSNR SSIM Param. (M) MACs (G) Time (s)
X MPRNet [51] 25.03 0.7081 20.13 10927 0.998
NAFNet [3] 25.06 0.7085 17.11 227 0.106
Uformer-B [47] 25.72 0.7334 50.88 2143 1.061
Stripformer [45] 25.93 0.7398 19.71 2397 1.367
FFTformer [21] 26.01 0.7481 16.56 1894 1.879
O DeepGyro [30] 23.78 0.6649 31.03 769 0.068
EggNet [16] 25.49 0.7266 6.34 1102 0.071
Ours 27.28 0.7803 16.31 262 0.130
Figure 7: Qualitative comparison on GyroBlur-Real. The red curves overlaid on the blurred images visualize the input gyro data.
@@
We first compare GyroDeblurNet with state-of-the-art single image deblurring methods that use gyro data  [30, 16] and those that do not [51, 3, 47, 45, 21]. We evaluate the performance of all the models on the test set of GyroBlur-Synth with gyro errors. The results are presented in Tab. 1. All the models in the table were trained with GyroBlur-Synth using their own hyperparameter settings.
@@
As shown in Tab. 1, GyroDeblurNet outperforms all the previous non-gyro-based methods by a large margin. Especially compared to Uformer-B [47], Stripformer [45] and FFTformer [21], which adopt transformers, our method requires a significantly smaller amount of computation, and a shorter computation time, while achieving more than 1 dB higher PSNR. Moreover, our method outperforms all the previous gyro-based methods by a large margin, while ours requires a comparable model size and a much smaller computational overhead.
@@
Qualitative comparisons on GyroBlur-Synth and GyroBlur-Real are shown in Fig. 6 and in Fig. 7, respectively. As shown in the figures, the previous non-gyro-based methods fail to restore sharp details due to severe blur. Moreover, despite the input gyro data, the previous gyro-based methods also fail to produce sharp details as they are not robust to gyro errors. On the other hand, our method successfully restores sharp images as our method can effectively exploit gyro data with errors.
Figure 8: Camera motion field visualization on a GyroBlur-Synth image and a GyroBlur-Real image. (a) camera motion field visualization on a GyroBlur-Synth image. (b) Our result of (a). (c) A GyroBlur-Real image. (d) Gyro errors. (e) Our result of (c).
Gyro errors
As discussed in Sec. 1, real-world gyro data cannot accurately represent the camera motion due to gyro noise and missing information such as rotational centers. Fig. 8 shows examples of such gyro errors in our datasets where camera motion fields are visualized as red curves. As shown in the figure, the camera motion information in the camera motion fields in both GyroBlur-Synth and GyroBlur-Real do not perfectly match the blur trajectories in the blurred images due to gyro errors. Nonetheless, GyroDeblurNet successfully restores sharp images for both synthetic and real-world input images.
Table 2: Results on GyroBlur-Synth with two models trained with noisy and noise-free gyro data.
    Gyro data Noisy Noise-free
  PSNR 27.28 24.94
  SSIM 0.7803 0.7112
Effect of training with noisy gyro data
Our training data includes gyro errors to generalize our model to real-world gyro errors. Here we demonstrate the effect of the synthetically added gyro errors. Fig. 9 shows that the model trained with noisy gyro data successfully removes blur from a real-world blurred image with gyro errors, while the one trained with noise-free gyro data fails. Tab. 2 presents a quantitative comparison on GyroBlur-Synth. As shown in the table, training with noisy gyro data substantially improves the deblurring quality by more than 2 dB, proving that including gyro errors in the training data is crucial for the training of gyro-based deblurring.
Figure 9: Effect of training with noisy gyro data. (a) Real-world blurred image. (b) Model trained with noisy gyro data. (c) Model trained with noise-free gyro data.
@@
Another source of gyro error is moving objects with different blur trajectories as discussed in Sec. 1. Fig. 10 shows a real-world example where the input blurry image has a moving object with different blur trajectory. Due to severe blur, the result of FFTformer [21], the state-of-the-art non-gyro-based method, still has remaining blur in both background and moving object. EggNet [16] also fails to deblur the image as it does not consider neither moving objects nor gyro errors. In contrast, our method successfully restores sharp details both on the moving object and the background thanks to its robustness to gyro errors.
Figure 10: Results of a real-world blurred image with a moving object.
Table 3: Effect of camera motion field hyperparameter $M$.
@@
We investigate the impact of the hyperparameter $M$ that determines the temporal resolution of the camera motion field. To verify the effect of $M$, we compare the deblurring performance of different values of $M$. As shown in Tab. 3, the deblurring quality improves gradually as $M$ increases. This highlights the importance of gyro data representation that can capture real-world intricate camera motions, and also explains the limited performance of previous gyro-based approaches [16, 30]. Fig. 11 also demonstrates the effect of increasing the number of vectors to handle complex real-world motion blur. However, increasing $M$ to 16 does not yield a significant performance gain while it doubles the memory consumption, indicating that $M=8$ is sufficient to represent the camera shakes in most of the images in the test set of GyroBlur-Synth.
Table 4: Ablation study on the components of GyroDeblurNet.
  Model PSNR SSIM
(a) Deblurring with no gyro data 24.90 0.6997
(b) Gyro refine. w/o image features 26.64 0.7584
(c) Deform. conv. using only gyro features 26.68 0.7588
(d) GyroDeblurNet 26.94 0.7667
(e) GyroDeblurNet + curriculum learning 27.28 0.7803
Ablation study
We conduct an ablation study to investigate the effect of individual components in our method. Specifically, we first study the impact of using gyro data. To this end, we build a variant of GyroDeblurNet that has no gyro module, and uses only image features for computing deformable convolution offsets in the gyro deblurring blocks. In our method, we compensate for gyro errors using image features in the gyro refinement and gyro deblurring blocks. Thus, we also examine the effect of using image features in these blocks using two variants of GyroDeblurNet. Specifically, we build one variant without the connections from the image deblurring module to the gyro module, and another variant that computes deformable convolution offsets using only gyro features. We trained all the aforementioned models and our full model without curriculum learning. Finally, to study the impact of the curriculum-learning-based training scheme, we also include our full model trained with curriculum learning.
Tab. 4 reports the comparison result. As shown in Tab. 4 (a), deblurring with no gyro data performs the worst as it cannot benefit from the blur cue provided by gyro data. The results in Tab. 4 (b) and (c) show that exploiting gyro data enhances deblurring quality, but their performances are still worse than our full model in (d) as they do not fully compensate for gyro errors. Finally, the curriculum learning (Tab. 4 (e)) substantially improves the deblurring quality as it helps train the model to fully utilize gyro data.

### v3

@@ 3.3 Training Strategy
To address the challenge of training a deblurring network with erroneous gyro data, we propose a curriculum learning-based training strategy. This approach, inspired by its success in handling noisy data in prior work [1, 38, 40, 43], aims to progressively guide the network in leveraging gyro information effectively despite inaccuracies.
Our training strategy initially trains the network with error-free gyro data, and gradually increases error in the training data. To this end, we first assume that, for each blurred image $B$, its corresponding error-free gyro data $G_{GT}$ is given as well as its ground-truth sharp image $S_{GT}$. From $G_{GT}$, we compute a noise-free camera motion field $\mathcal{V}_{clean}$ and a noisy camera motion field $\mathcal{V}_{noisy}$. We generate $\mathcal{V}_{noisy}$ by randomly perturbing the rotational center and adding noise to the gyro data $G_{GT}$. We then compute a blended camera motion field $\mathcal{V}_{\alpha}=(1-\alpha)\cdot\mathcal{V}_{clean}+\alpha\cdot\mathcal{V}_{noisy}$ where $\alpha$ is a blending parameter. We gradually increase $\alpha$ from 0 to 1 during training. For more details including the noisy camera motion field generation and the scheduling of $\alpha$, refer to the supplementary material.
@@ 4.1 GyroBlur-Synth
GyroBlur-Synth provides 14,600 and 640 synthetically blurred images of size $720\times 1280$ for the training and test sets. The dataset also provides ground-truth sharp images and the corresponding gyro data. To faithfully mimic real-world blurred images, the dataset covers various degradation sources, such as moving objects, noise, and saturation.
@@
To create synthetic blurred images with moving objects that have different blur trajectories, we used a simple method. We select a single instance from the COCO dataset [25] and assume it moves at a constant speed in a blurry image. We randomly choose its position and the direction and speed of motion. We then add the object to the warped sharp images before blending them during the blur synthesis process described earlier.
To reflect real-world noise and saturated pixels, we also adopt the blur synthesis pipeline of RSBlur [36]. Specifically, we synthesize blur in the linear sRGB space following the aforementioned process. Then, we synthesize saturated pixels, convert the image into the camera RAW space, synthesize noise, and convert the noisy RAW image to the linear sRGB image. We use the camera parameters of a Samsung Galaxy S22 ultra-wide camera, and the noise distribution estimated from it. We refer the readers to [36] for the detailed blur synthesis process.
To generate noisy camera motion fields, we add gyro errors with gyro data noise and randomly perturbed rotation centers. To obtain realistic gyro noise, we measured the gyro noise distribution from a stationary smartphone (Samsung Galaxy S22) placed on a table. Gyro noise is modeled as a Gaussian distribution for each rotation axis independently. For more details on the dataset, we refer the readers to the supplementary material.
Our synthetic dataset generation requires minimal device-specific input, brief gyro sequences, and simple camera calibration, significantly reducing manual data collection. We capture sufficient gyro data within minutes and obtain the necessary camera parameters for the RSBlur pipeline from a few calibration images. While we do not explicitly model the gyro’s position relative to the camera, this is implicitly addressed as gyro errors such as rotational center error. Our experiments demonstrate effective blur removal regardless of these errors.
Method Models GyroBlur-Synth GyroBlur-Real-S Param. (M) MACs (G) Time (s)
PSNR $\uparrow$ SSIM $\uparrow$ NIQE $\downarrow$ TOPIQ $\uparrow$
Single-image deblurring MPRNet [52] 25.03 0.7081 5.25 0.397 20.13 10927 0.998
NAFNet [4] 25.06 0.7085 5.27 0.409 17.11 227 0.106
AdaRevD-B (NAFNet) [27] 25.61 0.7281 5.18 0.425 43.81 2658 1.661
Uformer-B [48] 25.72 0.7334 4.80 0.431 50.88 2143 1.061
Stripformer [45] 25.93 0.7398 4.71 0.456 19.71 2397 1.367
FFTformer [19] 26.01 0.7481 4.98 0.434 16.56 1894 1.879
Non-blind Vasu et al. [47] 19.63 0.4526 6.13 0.317 10.41 - 160.176
deblurring Nan et al. [32] 22.22 0.5313 5.81 0.348 26.16 41170 27.069
Gyro-based deblurring DeepGyro [30] 23.78 0.6649 5.64 0.381 31.03 769 0.068
EggNet [15] 25.49 0.7266 5.18 0.413 6.34 1102 0.071
INformer [34] 25.11 0.7103 5.29 0.408 24.88 1315 0.464
Ours 27.28 0.7803 4.47 0.548 16.31 262 0.130
Table 1: Quantitative comparison on the GyroBlur-Synth and GyroBlur-Real-S. For the GyroBlur-Real-S, we use non-reference metrics (NIQE [28] and TOPIQ [2] trained on KonIQ-10k dataset [12]) for evaluation. Inference times were measured using a $720\times 1280$ image. For models utilizing gyro data, inference times include both gyro data embedding construction and the deblurring process.
@@ 4.2 GyroBlur-Real
GyroBlur-Real consists of two subsets: GyroBlur-Real-S and GyroBlur-Real-H. GyroBlur-Real-S was collected using a Samsung Galaxy S22, which was also used for constructing the GyroBlur-Synth dataset. We use GyroBlur-Real-S as the primary dataset to evaluate the performance of our approach on real-world images. On the other hand, GyroBlur-Real-H was collected using a Huawei P30 Pro, and is utilized for evaluating the generalization ability of our approach. GyroBlur-Real-S provides 100 real-world blurry images of size $4080\times 3060$ along with their corresponding gyro data. GyroBlur-Real-H provides 17 real-world blurry images of size $5120\times 3840$ and their gyro data. The images in both datasets are provided in DNG and JPEG.
@@ 5 Experiments
We implemented our network using PyTorch. We trained our network on randomly cropped $256\times 256$ image patches from the GyroBlur-Synth dataset for 300 epochs with a batch size of 16. We used the Adam optimizer [18] with $\beta_{1}=0.9$ and $\beta_{2}=0.999$. The learning rate was initially set to 0.0001 and reduced to 1e-7 using the cosine annealing scheduler [26]. We used the PSNR loss [3, 4] to train our model. We measured the computation times of all the models on a GeForce RTX 3090 GPU.
We evaluate the performance of our method using the test set of GyroBlur-Synth and GyroBlur-Real. Additional experiments and examples including visualization of the effect of gyro refinement and an analysis on the robustness to gyro errors are included in the supplementary material.
Figure 6: Qualitative comparison on GyroBlur-Real-S. The red curves overlaid on the blurred images visualize the input gyro data.
@@
We compare GyroDeblurNet with state-of-the-art single image deblurring methods that use gyro data [30, 15, 34] and those that do not [52, 4, 48, 45, 19, 27]. We also compare GyroDeblurNet with non-blind deconvolution methods that are designed to handle kernel error [32, 47]. Since these methods assume uniform blur kernels, we extended them to handle spatially-varying blur kernels computed from gyro data by using patch-wise blur kernels [9]. For more details about our extensions of the non-blind deconvolution methods, we refer the readers to the supplementary material. We train all models with the training set of GyroBlur-Synth and evaluate the performance of all the models on the test set of GyroBlur-Synth and GyroBlur-Real-S.
@@
The results are presented in Tab. 1. As shown in the table, GyroDeblurNet significantly outperforms all previous non-gyro-based methods. Specifically, compared to transformer-based approaches [48, 45, 19], our method requires substantially less computation and shorter computation time, while achieving more than 1 dB higher PSNR. Furthermore, our method surpasses all previous gyro-based methods by a large margin, despite requiring a comparable model size and much smaller computational overhead. The non-blind deblurring methods designed to handle blur kernel errors also perform significantly worse than our method due to large gyro errors. Nonetheless, our approach demonstrates superior performance even in the presence of large gyro errors.
@@
Qualitative comparisons on GyroBlur-Synth and GyroBlur-Real-S are shown in Fig. 5 and in Fig. 6, respectively. As shown in the figures, the previous non-gyro-based methods fail to restore sharp details due to severe blur. Moreover, despite the input gyro data, the previous gyro-based methods also fail to produce sharp details as they are not robust to gyro errors. The non-blind deblurring methods that are designed to handle kernel errors show significant ringing artifacts due to large errors in the gyro data. On the other hand, our method successfully restores sharp images as our method can effectively exploit erroneous gyro data.
Model GyroBlur-Synth GyroBlur-Real-S
PSNR $\uparrow$ / SSIM $\uparrow$ NIQE $\downarrow$ / TOPIQ $\uparrow$
(a) Deblurring w/ no gyro data 24.90 / 0.6997 5.23 / 0.391
(b) Training w/ error-free gyro data 24.94 / 0.7112 5.27 / 0.396
(c) Deblurring w/ gyro data w/o any refine. 25.47 / 0.7128 5.01 / 0.418
(d) Gyro refine. w/o image features 26.17 / 0.7465 4.75 / 0.447
(e) Deform. conv. using only gyro features 26.32 / 0.7542 4.76 / 0.460
(f) GyroDeblurNet 26.94 / 0.7667 4.61 / 0.511
(g) GyroDeblurNet + curriculum learning 27.28 / 0.7803 4.47 / 0.548
Table 2: Ablation study on the components of GyroDeblurNet. For the GyroBlur-Real-S, we use non-reference metrics (NIQE [28] and TOPIQ [2] trained on KonIQ-10K dataset [12]) for evaluation.
Figure 7: Qualitative results of the ablation study on a real-world blurry image.
Ablation study
We conducted an ablation study to investigate the effects of individual components in our method. To this end, we built several variants of GyroDeblurNet and compared their performances. Both quantitative and qualitative results are reported in Tab. 2 and Fig. 7, respectively.
Tab. 2-(a) corresponds to a model with no gyro modules, using only image features in the gyro deblurring blocks. Tab. 2-(b) represents our full model, but trained with error-free gyro data. Tab. 2-(c) is a model that is trained with erroneous gyro data but does not refine gyro features using image features. Specifically, in this model, gyro features are not concatenated with image features in both gyro refinement and gyro deblurring blocks. Tab. 2-(f) is our full model trained without curriculum learning, while Tab. 2-(g) is our full model trained with curriculum learning.
As shown in Tab. 2-(a), deblurring with no gyro data performs the worst, as it cannot benefit from the blur cue provided by gyro data. The model in (b) achieves only marginal improvement over (a), indicating the importance of reflecting gyro errors in training data. The results in (c), (d), (e) and (f) demonstrate the necessity of training with erroneous gyro data and gyro refinement using image features. Finally, curriculum learning in (g) further improves deblurring quality, as it helps the model to fully utilize erroneous gyro data.
Fig. 7 also highlights the impact of each component in our method. The results in Fig. 7-(b) and (c) reveal that training with error-free gyro data yields no performance gain over training without gyro data since error-free gyro data cannot reflect real-world gyro errors. In contrast, Fig. 7-(d) shows that training with erroneous gyro data aids in handling these errors but remains suboptimal. Gyro refinement using image features in Fig. 7-(e) significantly improves image quality, while Fig. 7-(f) demonstrates how curriculum learning further enhances error handling.
Figure 8: Results of a real-world blurred image with a moving object. (a) Blurry image and its background region and foreground object. (b) Result of FFTformer. (c) Result of GyroDeblurNet.
@@
Another source of gyro error is moving objects with different blur trajectories. Fig. 8 shows a real-world example where the input blurry image has a moving object with a different blur trajectory. Due to severe blur, the result of FFTformer [19], the state-of-the-art non-gyro-based method, still has remaining blur in the background. In contrast, our method successfully restores sharp details on the background and shows comparable quality on the moving object thanks to its robustness to gyro errors.
@@
Table 3: Effect of camera motion field hyperparameter $M$.
@@
We investigate the impact of the hyperparameter $M$ that determines the temporal resolution of the camera motion field. As shown in Tab. 3, the deblurring quality improves gradually as $M$ increases. This highlights the importance of gyro data representation that can capture real-world intricate camera motions, and also explains the limited performance of previous gyro-based approaches [15, 30]. Fig. 9 also demonstrates the effect of increasing the number of vectors to handle complex real-world motion blur. However, increasing $M$ to 16 does not yield a significant performance gain while it doubles the memory consumption, indicating that $M=8$ is sufficient to represent the camera shakes in most of the images in the test set of GyroBlur-Synth.
Generalization to other devices
We test GyroDeblurNet’s generalization ability by comparing it with top deblurring methods from Tab. 1 on GyroBlur-Real-H. All methods were trained on GyroBlur-Synth. The results in Tab. 4 show our approach outperforms the others, even on a different device. Fig. 10 also demonstrates that our approach successfully removes blur from an image captured by a different device, highlighting its generalization ability.
Figure 10: Blurry image taken with a Huawei P30 Pro device and its deblurred result. (a) A blurry image and its gyro errors. (b) Deblurred result of (a).
  Model Stripformer [45] FFTformer [19] EggNet [15] Ours
NIQE $\downarrow$ 5.47 5.22 5.71 4.94
TOPIQ $\uparrow$ 0.451 0.459 0.432 0.521
Table 4: Quantitative evaluation on GyroBlur-Real-H using non-reference metrics [28, 2].